OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0).

OpenStudy (anonymous):

First Find the Slope

OpenStudy (anonymous):

\[m=\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

Name any points by (x1, y1) and (x2, y2) and use the distance formula. Distance= radical ( (x2-x1)^2 + (y2-y1) All of them are under the radical.

OpenStudy (anonymous):

Nw use pont Slope Form... y-y1 = m(x-x1)

OpenStudy (anonymous):

Nw rearrange the Equation into.... y = mx + c

OpenStudy (anonymous):

Hope this helps..)

OpenStudy (anonymous):

so its Y+6=\[\frac{ 1 }{ 2 }\] (x-4)

OpenStudy (anonymous):

Yup..)

OpenStudy (anonymous):

y+ 6 = 1/2 ( x-4) nw write this in the form y=mx+ c